# Bridging Speech Emotion Recognition and Personality: Dataset and Temporal Interaction Condition Network

### v1

@@ Abstract
This study investigates the interaction between personality traits and emotional expression, exploring how personality information can improve speech emotion recognition (SER). We collected personality annotation for the IEMOCAP dataset, and the statistical analysis identified significant correlations between personality traits and emotional expressions. To extract finegrained personality features, we propose a temporal interaction condition network (TICN), in which personality features are integrated with Hubert-based acoustic features for SER. Experiments show that incorporating ground-truth personality traits significantly enhances valence recognition, improving the concordance correlation coefficient (CCC) from 0.698 to 0.785 compared to the baseline without personality information. For practical applications in dialogue systems where personality information about the user is unavailable, we develop a front-end module of automatic personality recognition. Using these automatically predicted traits as inputs to our proposed TICN model, we achieve a CCC of 0.776 for valence recognition, representing an 11.17% relative improvement over the baseline. These findings confirm the effectiveness of personality-aware SER and provide a solid foundation for further exploration in personality-aware speech processing applications.
Index Terms:
Speech emotion recognition, Big Five personality traits, human computer interaction.
@@ I Introduction
Speech emotion recognition (SER) is widely known as a vital component of natural human–computer interaction [1, 2]. Emotional information in speech reveals not only the immediate affective state of the speaker but also provides insight into how we think, feel, and behave [3]. By enabling intelligent systems to detect and respond to the emotional state of the user, SER makes interactions more intuitive and improves the overall user experience. Applications of SER range from empathetic speech assistants [4] and personalized systems for emotion-aware healthcare and automotive interfaces [5], underscoring its significance within the broader field of affective computing.
@@
Moreover, how we express and manage emotions is significantly influenced by our personality traits [6]. These traits shape cognitive processes, reactions, and communication styles, and thus play a key role in social interactions and emotional expression. Among various personality models [7, 8], [9], the Big Five traits [10, 11] is the most widely adopted framework in psychological research [12, 13]. It defines five traits: openness (OP), conscientiousness (CO), extraversion (EX), agreeableness (AG), and neuroticism (NE). Each trait represents a distinct dimension of personality and offers a practical solution to quantify individual differences [14]. Previous studies have shown that these traits can substantially influence emotion perception, regulation, and expression [15, 16]. For instance, individuals high in EX tend to express their emotions more openly and with higher arousal, whereas those high in NE are often more sensitive to negative emotions and exhibit greater emotional instability.
@@
In this work, we annotate the well-known Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [17] with Big Five traits. This personality-annotated IEMOCAP (PA-IEMOCAP) dataset, to our knowledge, is the first dataset of this type for speech, enables direct investigation of the interplay between personality and emotion expression  [^1] The GitHub repository containing the data will be made publicly available when the paper is accepted.. We first examine the correlation between each Big Five trait and the emotional dimensions of valence and arousal. Additionally, we incorporate a straightforward linear layer to project personality traits as feature embedding, which is integrated with acoustic features to improve the performance of SER.
@@
In practical applications, explicit personality information about users is unavailable [18]. In this work, we explore effective models that incorporate automatically predicted personality traits for improving SER. Our approach begins by establishing a baseline for personality recognition (PR) at the utterance level. We then conduct multi-task learning experiments, jointly training PR and SER to explore their mutual influence within a shared feature extractor. Given that personality traits remain consistent throughout a conversation, we further provide a conversational-level PR baseline. These experiments reveal the potential of real dialogue systems in personalizing response styles according to the personality of users.
@@
Additionally, the predicted personality traits can serve as input of condition network, replacing ground-truth personality labels and thereby enable dialogue system better interpret and recognize emotional states about user. To better incorporate this predicted personality information for SER, we introduce a temporal interaction condition network (TICN). Although personality remains consistent over time, its impact on emotion expressions can vary across different speech segments. The proposed TICN is designed to capture this temporal variation, extracting finegrained personality features that enhance SER performance. Instead of combining features by concatenation, we incorporate cross attention mechanism [19], allowing effective interaction between personality and acoustic features.
@@
  • Personality annotated IEMOCAP dataset: We augment the widely-used IEMOCAP dataset with Big Five personality trait annotations, providing the first speech dataset that includes both emotion and personality labels.
@@ II-A Attribute Information for SER
SER has attracted increasing attention in recent years. Prior studies have shown that including additional speaker attributes such as gender [20], vocal characteristics [21], and linguistic content [22] can improve SER model performance. Researchers introduce multi-task learning to integrate these speaker attributes into SER models [23, 24], [25]. By simultaneously optimizing multiple related tasks, multi-task learning enables models to learn shared features and better utilize complementary information for emotion recognition [26, 27]. This approach has shown promising performance in SER [28, 29] benefiting from the extra information provided by emotion-related tasks.
Regarding the relationship between linguistic content and emotions, Cai et al. [22] proposed a multi-task learning method combining ASR and SER using the wav2vec 2.0 feature extractor. Joint training for ASR and SER allowed the model to leverage linguistic information effectively, leading to state-of-the-art performance on the IEMOCAP dataset. An ablation study further confirmed that carefully choosing the weight parameters for ASR significantly improved results. Additionally, speaker attributes have been widely studied: different speakers express emotions differently, affecting both acoustic and linguistic content for the same emotions. Thus, speaker-dependent SER systems usually outperform speaker-independent ones [30]. As a broad category of speakers, gender information can also improve performance in speaker-independent SER [20, 31]. More recently, Sharma et al. [28] proposed a multi-task learning approach using pretrained self-supervised learning feature extractor. Their study showed further advantages of multi-task learning by integrating additional tasks like language classification and regression tasks related to pitch and energy. These additional tasks improved SER performance significantly across 25 datasets covering 13 languages and 7 emotion categories, demonstrating the effectiveness of multi-task learning in enhancing model generalization, especially in multilingual settings. Despite the known relationship between personality traits and emotional expressions, personality remains an underexplored auxiliary attribute in SER.
@@ II-B Joint Analysis of Personality and Emotion in Speech
Personality traits and emotional states are closely connected, both influencing how people express themselves [32, 33]. Understanding how these two aspects interact can lead to better performance in SER tasks. Recently, researchers have started exploring joint analysis frameworks to capture the interaction between personality and emotion, even though the lack of datasets annotated with both attributes remains a major challenge. Zhang et al. [34] proposed PersEmoN, a deep multi-task learning framework for jointly analyzing personality and emotion. Because no publicly available corpus contained both emotion and personality annotations, they used two separate datasets (one labeled for emotion and another for personality) and incorporated an adversarial learning method to mitigate the mismatch of different datasets. Their results highlighted an inherent relationship between personality and emotion: personality shows stable impact on how emotions are expressed and managed. For example, they observed that EX corresponds to more intense emotions, while NE correlates with more frequent negative expression. Similarly, Li et al. [35] explored the personality-emotion connection in a transfer learning manner. They trained a wav2vec2-based model [36] for SER and then finetuned it for PR, finding that emotional features, particularly the arousal dimension, significantly improve PR performance. These studies underscore the effectiveness of incorporating personality information into SER.
@@ III Personality Annotation for IEMOCAP dataset
We conducted personality annotations for the widely-used Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset to explore relationships between personality traits and emotional expressions. The IEMOCAP dataset [17] is a benchmark dataset extensively used in emotion recognition and sentiment analysis research. It comprises approximately 12 hours of multimodal data collected from 10 professional actors (5 males and 5 females), including high-quality audio recordings, video, detailed facial motion capture data, and textual transcriptions. Data collection involved five dyadic sessions, each including one male and one female actor performing scripted conversations and engaging in improvisational scenarios designed to elicit specific emotional expressions. The audio was captured using two microphones at a 48 kHz sampling rate and subsequently downsampled to 16 kHz to align with the common audio processing standard. Below, we introduce key aspects of our annotation method, consistency evaluation, and correlation analysis.
@@ III-A Labeling Procedure
In the IEMOCAP dataset, professional actors perform diverse roles in each conversation. We presume they play different characters in each conversation, and we annotated the Big Five personality traits at the conversation level, resulting in 302 personality profiles for 151 dyadic interactions. The annotation was performed by six independent raters recruited via Amazon Mechanical Turk, each tasked with evaluating the personality traits of individual speakers based on both the audio recordings and transcribed text of the conversations. The Ten Item Personality Measure (TIPI) [37], a widely validated instrument for assessing the Big Five personality dimensions (OP, CO, EX, AG, and NE), was employed, with raters providing scores on a 7-point Likert scale (ranging from 1 = “strongly disagree” to 7 = “strongly agree”) for each trait. To ensure the quality and consistency of the annotations, we implemented a rigorous data-cleaning procedure. Specifically, we excluded ratings deemed contradictory, such as instances where a speaker was simultaneously rated as “critical, quarrelsome” and “sympathetic, warm” for the Agreeableness trait, as such inconsistencies could undermine the reliability of the personality profiles. Furthermore, to mitigate the impact of outliers and enhance robustness, we removed the single rating furthest from the median score for each trait, subsequently computing the average of the remaining five ratings to generate a final trait score.
@@
TABLE I: Inter-annotator agreement analysis for Big Five personality trait annotations. We report both ICC(2,k) and Fleiss’ kappa values to evaluate the consistency among annotators.
Personality traits ICC Fleiss’s kappa
OP 0.89 0.49
CO 0.89 0.53
EX 0.89 0.47
AG 0.97 0.83
NE 0.95 0.71
Average 0.92 0.61
@@
TABLE II: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence. Values marked with indicate a strong correlation ($>|0.5|$).
Personality traits Valence Arousal
OP $0.53^{\dagger}$ $0.09$
CO $0.32$ $-0.22$
EX $0.35$ $0.32$
AG $0.51^{\dagger}$ $-0.11$
NE $-0.45$ $0.19$
III-B Consistency Assessment
@@
To evaluate the reliability of personality annotations, we introduce two metrics: intraclass correlation coefficient (ICC) [38] for the collected continuous annotation (1-7 scale) and Fleiss’ kappa [39] for binarized (median-split) traits. As shown in Table I, ICC values demonstrate excellent agreement across all traits, ranging from 0.89 to 0.97 for the entire database, indicating strong reliability in the continuous personality assessments. We also ensure annotation reliability by converting the continuous personality trait scores into binary labels, maintaining robustness in simple personality classifications (e.g., extrovert vs. introvert). We observe that the average $\kappa$ across all annotations was 0.61. Among individual traits, AG and NE exhibited the highest agreement ($\kappa$=0.83 and $\kappa$=0.71, respectively), while EX had the lowest ($\kappa$=0.47). These results confirm the reliability and consistency of our personality annotations. The resultant annotated dataset is thereafter referred to as PA-IEMOCAP.
III-C Correlation of Big Five Traits and Emotions
@@
We introduce Pearson correlation coefficients (PCC) [40] to assess the correlation between personality traits and the emotion dimensions of valence and arousal (Table II). Statistically significant correlations ($p<0.05$) were observed, notably OP showing strong positive correlations with valence (PCC = 0.53), which implies that individuals scoring high in OP tend to exhibit more positive emotions. NE, conversely, shows a pronounced negative correlation with valence (PCC = -0.45), indicating individuals with higher NE levels express more negative emotions. AG and CO also present noticeable associations with valence. Regarding arousal, EX has a positive correlation (PCC = 0.32), suggesting extroverts typically display heightened emotional energy. These results provide the first quantitative analysis of personality and emotion expression.

### v2

@@ Abstract
This study investigates the interaction between personality traits and emotion expression, exploring how personality information can improve speech emotion recognition (SER). We collect the personality annotation for the IEMOCAP dataset, making it the first speech dataset that contains both emotion and personality annotations (PA-IEMOCAP), and enabling direct integration of personality traits into SER. Statistical analysis on this dataset identified significant correlations between personality traits and emotional expressions. To extract finegrained personality features, we propose a temporal interaction condition network (TICN), in which personality features are integrated with HuBERT-based acoustic features for SER. Experiments show that incorporating ground-truth personality traits significantly enhances valence recognition, improving the concordance correlation coefficient (CCC) from 0.698 to 0.785 compared to the baseline without personality information. For practical applications in dialogue systems where personality information about the user is unavailable, we develop a front-end module of automatic personality recognition. Using these automatically predicted traits as inputs to our proposed TICN model, we achieve a CCC of 0.776 for valence recognition, representing an 11.17% relative improvement over the baseline. These findings confirm the effectiveness of personality-aware SER and provide a solid foundation for further exploration in personality-aware speech processing applications.
@@ I Introduction
Speech emotion recognition (SER) is widely known as a vital component of natural human–computer interaction [cowie2001emotion, ramakrishnan2013speech]. Emotional information in speech reveals not only the immediate affective state of the speaker but also provides insight into how we think, feel, and behave [vernon2006thinking]. By enabling intelligent systems to detect and respond to the emotional state of the user, SER makes interactions more intuitive and improves the overall user experience. Applications of SER range from empathetic speech assistants [chatterjee2021real] and personalized systems for emotion-aware healthcare and automotive interfaces [dhuheir2021emotion], underscoring its significance within the broader field of affective computing.
@@
Moreover, how we express and manage emotions is significantly influenced by our personality traits [stemmler2010personality]. These traits shape cognitive processes, reactions, and communication styles, and thus play a key role in social interactions and emotional expression. Among various personality models [cattell1992handbook, ashton2007empirical], [myers2010gifts], the Big Five traits [costa1999five, costa2008revised] is the most widely adopted framework in psychological research [azucar2018predicting, simha2020big]. It defines five traits: openness (OP), conscientiousness (CO), extraversion (EX), agreeableness (AG), and neuroticism (NE). Each trait represents a distinct dimension of personality and offers a practical solution to quantify individual differences [langston1997beliefs]. Previous studies have shown that these traits can substantially influence emotion perception, regulation, and expression [deniz2011investigation, curtis2015relationship]. For instance, individuals high in EX tend to express their emotions more openly and with higher arousal, whereas those high in NE are often more sensitive to negative emotions and exhibit greater emotional instability.
@@
In this work, we annotate the well-known Interactive Emotional Dyadic Motion Capture (IEMOCAP) dataset [busso2008iemocap] with Big Five traits. This personality-annotated IEMOCAP (PA-IEMOCAP) dataset, to our knowledge, is the first dataset of this type for speech, enables direct investigation of the interaction between personality and emotion expression [^1] Data available at https://github.com/Kyoto-University-Speech-and-Audio/PA-IEMOCAP. We first examine the correlation between each Big Five trait and the emotional dimensions of valence and arousal. Additionally, we incorporate a straightforward linear layer to project personality traits as feature embedding, which is integrated with acoustic features to improve the performance of SER.
@@
In practical applications, explicit personality information about users is unavailable [wu2017implicit]. In this work, we explore effective models that incorporate automatically predicted personality traits for improving SER. Our approach begins by establishing a baseline for personality recognition (PR) at the utterance level. We then conduct multi-task learning experiments, jointly training PR and SER to explore their mutual influence within a shared feature extractor. Given that personality traits remain consistent throughout a conversation, we further provide a conversational-level PR baseline. These experiments reveal the potential of real dialogue systems in personalizing response according to the personality of users.
@@
Additionally, the predicted personality traits can serve as input of condition network, replacing ground-truth personality labels and thereby enable dialogue system better interpret and recognize emotional states about user. To better incorporate this predicted personality information for SER, we introduce a temporal interaction condition network (TICN). Although personality remains consistent over time, its impact on emotion expressions can vary across different speech segments. The proposed TICN is designed to capture this temporal variation, extracting finegrained personality features that enhance SER performance. Instead of combining features by concatenation, we incorporate cross attention mechanism [vaswani2017attention], allowing effective interaction between personality and acoustic features.
@@
  • PA-IEMOCAP dataset: We augment the widely-used IEMOCAP dataset with Big Five personality trait annotations, providing the first speech dataset that includes both emotion and personality labels.
@@ II-A Attribute Information for SER
SER has attracted increasing attention in recent years [khan2025joint]. Prior studies have shown that including additional speaker attributes such as gender [li2019improved], vocal characteristics [triantafyllopoulos2021deep], and linguistic content [cai2021speech] can improve SER model performance. Researchers introduce multi-task learning to integrate these speaker attributes into SER models [collobert2008unified, standley2020tasks], [zhang2021survey]. By simultaneously optimizing multiple related tasks, multi-task learning enables models to learn shared features and better utilize complementary information for emotion recognition [zhang2017cross, sener2018multi]. This approach has shown promising performance in SER [sharma2022multi, latif2022multitask] benefiting from the extra information provided by emotion-related tasks.
In multimodal emotion recognition, combining speech and text consistently outperforms speech-only systems, as linguistic information provides complementary semantic cues. Khan et al. [khan2025memocmt] introduced a cross-modal transformer that integrates HuBERT-based speech features with BERT-based text features, achieving state-of-the-art results on benchmark datasets. These findings highlight the importance of jointly modeling acoustic and linguistic cues for emotion recognition. When explicit transcripts are unavailable, automatic speech recognition (ASR) can serve as an auxiliary task to implicitly capture linguistic information from speech. Cai et al. [cai2021speech] proposed a multi-task learning framework combining ASR and SER with the wav2vec 2.0 feature extractor, where joint training effectively leveraged linguistic information and achieved state-of-the-art performance on IEMOCAP. Their ablation study further demonstrated that carefully weighting the ASR loss was critical to performance gains. Additionally, speaker attributes have been widely studied: different speakers express emotions differently, affecting both acoustic and linguistic content for the same emotions. Thus, speaker-dependent SER systems usually outperform speaker-independent ones [gat2022speaker]. As a broad category of speakers, gender information can also improve performance in speaker-independent SER [li2019improved, pappagari2020x]. More recently, Sharma et al. [sharma2022multi] proposed a multi-task learning approach using pretrained self-supervised learning feature extractor. Their study showed further advantages of multi-task learning by integrating additional tasks like language classification and regression tasks related to pitch and energy. These additional tasks improved SER performance significantly across 25 datasets covering 13 languages and 7 emotion categories, demonstrating the effectiveness of multi-task learning in enhancing model generalization, especially in multilingual settings. Despite the known relationship between personality traits and emotional expressions, personality remains an underexplored auxiliary attribute in SER.
@@ II-B Joint Analysis of Personality and Emotion in Speech
Personality traits and emotional states are closely connected, both influencing how people express themselves [sagha2017effect, guidi2019analysis]. Understanding how these two aspects interact can lead to better performance in SER tasks. Recently, researchers have started exploring joint analysis frameworks to capture the interaction between personality and emotion, even though the lack of datasets annotated with both attributes remains a major challenge. Zhang et al. [zhang2019persemon] proposed PersEmoN, a deep multi-task learning framework for jointly analyzing personality and emotion. Because no publicly available corpus contained both emotion and personality annotations, they used two separate datasets (one labeled for emotion and another for personality) and incorporated an adversarial learning method to mitigate the mismatch of different datasets. Their results highlighted an inherent relationship between personality and emotion: personality shows stable impact on how emotions are expressed and managed. For example, they observed that EX corresponds to more intense emotions, while NE correlates with more frequent negative expression. Similarly, Li et al. [li2023transfer] explored the personality-emotion connection in a transfer learning manner. They trained a wav2vec2-based model [baevski2020wav2vec] for SER and then finetuned it for PR, finding that emotional features, particularly the arousal dimension, significantly improve PR performance. These studies underscore the effectiveness of incorporating personality information into SER.
@@ III Personality Annotation for IEMOCAP dataset
We conducted personality annotations for the widely-used IEMOCAP dataset to explore relationships between personality traits and emotional expressions. The IEMOCAP dataset [busso2008iemocap] is a benchmark dataset extensively used in emotion recognition and sentiment analysis research. It comprises approximately 12 hours of multimodal data collected from 10 professional actors (5 males and 5 females), including high-quality audio recordings, video, detailed facial motion capture data, and textual transcriptions. Data collection involved five dyadic sessions, each including one male and one female actor performing scripted conversations and engaging in improvisational scenarios designed to elicit specific emotional expressions. The audio was captured using two microphones at a 48 kHz sampling rate and subsequently downsampled to 16 kHz to align with the common audio processing standard. Below, we introduce key aspects of our annotation method, consistency evaluation, and correlation analysis.
@@ III-A Labeling Procedure
To elicit genuine emotional expressions and provide a more natural representation of affective behavior, the IEMOCAP dataset employs professional actors who perform diverse roles across conversations. We assume that the actors portrayed different characters in a manner closely approximating real-life behavior, and we therefore annotated the Big Five personality traits at the conversation level. These annotations reflected the enacted personalities in each conversation, rather than the actual personalities of the speakers, resulting in 302 personality profiles derived from 151 dyadic interactions. The annotation was performed by six independent raters recruited via Amazon Mechanical Turk, each tasked with evaluating the personality traits of individual speakers based on both the audio recordings and transcribed text of the conversations. To ensure annotation accuracy, annotators were recruited from the same sociocultural background as the original IEMOCAP speakers, i.e., native speakers of American English. The Ten Item Personality Measure (TIPI) [gosling2003very], a widely validated instrument for assessing the Big Five personality dimensions (OP, CO, EX, AG, and NE), was employed, with raters providing scores on a 7-point Likert scale (ranging from 1 = “strongly disagree” to 7 = “strongly agree”) for each trait. To ensure the quality and consistency of the annotations, we implemented a rigorous data-cleaning procedure. Specifically, we excluded ratings deemed contradictory, such as instances where a speaker was simultaneously rated as “critical, quarrelsome” and “sympathetic, warm” for the AG trait, as such inconsistencies could undermine the reliability of the personality profiles. All problematic annotations were discarded, and new raters were recruited to replace them, ensuring that each conversation ultimately received six valid and independent ratings in total. The final trait score was computed as the average of the six ratings, without applying any further normalization.
@@
TABLE I: Inter-annotator agreement analysis for Big Five personality trait annotations. We report ICC(2,k) with 95% confidence intervals (CI 95%) and Fleiss’ kappa to evaluate the consistency among annotators. ICC(3,k) is additionally reported as a sensitivity check.
Personality traits ICC(2,k), CI 95% ICC(3,k) Fleiss’ kappa
OP 0.89, [0.87, 0.91] 0.90 0.49
CO 0.89, [0.87, 0.91] 0.89 0.53
EX 0.89, [0.87, 0.91] 0.89 0.47
AG 0.97, [0.96, 0.97] 0.97 0.83
NE 0.95, [0.94, 0.96] 0.95 0.71
Average 0.92 0.92 0.61
@@
TABLE II: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence at the utterance level. Subgroup analyses are reported for male and female speakers. Values marked with indicate a strong correlation ($|PCC|>0.5$).
Overall ($N=10039$) Male ($N=5239$) Female ($N=4800$)
Personality Valence Arousal Valence Arousal Valence Arousal
[…]
CO $0.33$ 0.21-0.21 $0.28$ 0.19-0.19 $0.38$ 0.24-0.24
EX $0.35$ $0.32$ $0.35$ $0.29$ $0.36$ $0.34$
[…]
TABLE III: Pearson correlation coefficients (PCC) between the Big Five personality traits and arousal and valence at the conversation level. Subgroup analyses are reported for male and female speakers. Values marked with indicate a strong correlation ($|PCC|>0.5$).
Overall ($N=302$) Male ($N=151$) Female ($N=151$)
Personality Valence Arousal Valence Arousal Valence Arousal
[…]
CO $0.41$ 0.41-0.41 $0.33$ 0.36-0.36 $0.49$ 0.48-0.48
[…]
To evaluate the reliability of personality annotations, we employed two metrics: the intraclass correlation coefficient (ICC) [fisher1970statistical] for the continuous ratings (1–7 scale) and Fleiss’ kappa [fleiss2013statistical] for the binarized traits obtained via median split. Before the consistency analysis, we excluded the rating that deviated the most from the median for each trait. As shown in Table I, inter-annotator agreement was excellent across all traits, with an average ICC(2,k) of 0.92. After data cleaning, the ICC(2,k) values ranged from 0.89 to 0.97, with 95% confidence intervals consistently above 0.85. As a sensitivity check, we additionally report ICC(3,k), which yielded nearly identical values, confirming the robustness of the reliability estimates. Without trimming the ratings furthest from the median score (mentioned in Section 3(A)), the average ICC(2,k) was 0.83, which still indicates strong reliability.
@@
We also ensure annotation reliability by converting the continuous personality trait scores into binary labels using a median split, thereby maintaining robustness in simplified personality classifications (e.g., extrovert vs. introvert). The average Fleiss’ $\kappa$ across all traits was 0.61; without trimming extreme ratings, the average $\kappa$ decreased to 0.39. Among individual traits, AG and NE exhibited the highest agreement ($\kappa$=0.83 and $\kappa$=0.71, respectively), while EX had the lowest ($\kappa$=0.47), likely because the conversational context did not clearly reflect introversion–extroversion, making it harder to annotate. These results confirm the reliability and consistency of our personality annotations. The resultant annotated dataset is thereafter referred to as PA-IEMOCAP.
III-B Correlation of Big Five Traits and Emotions
@@
We introduce Pearson correlation coefficients (PCC) [cohen2009pearson] between the Big Five traits and the two emotion dimensions, valence and arousal, at utterance level (Table II). Statistically significant correlations ($p<0.05$) were observed: OP shows strong positive correlations with valence (PCC = 0.53), which implies that individuals scoring high in OP tend to exhibit more positive emotions. NE shows a pronounced negative correlation with valence (PCC = -0.45), indicating individuals with higher NE levels express more negative emotions. AG and CO also exhibit noticeable associations with valence. Regarding arousal, EX has a positive correlation (PCC = 0.32), suggesting extroverts typically display heightened emotional energy.
Because personality labels are constant within a conversation and thus repeated across a speaker’s utterances, treating utterances as independent may violate the independence assumption. We therefore averaged valence and arousal for each speaker in a conversation and computed PCCs across conversations (Table III), thereby capturing the association between personality and overall emotion expression. Compared with the utterance-level analysis, correlations show modest increases overall; in particular, the association between Extraversion and arousal increases to (PCC=0.61). We hypothesize that aggregating at the speaker level reduces within-speaker noise, thereby yielding better correlations.
In addition, we conducted subgroup analyses by gender. The results for male and female speakers show broadly consistent patterns with the overall population, and no substantial differences between the genders were observed. For instance, the OP–valence and EX–arousal associations remained strong across both groups, while NE consistently exhibited negative correlations with valence.